# NüshuRescue: Revitalization of the endangered Nüshu Language with AI

### v2

@@ Abstract
The preservation and revitalization of endangered and extinct languages is a meaningful endeavor, conserving cultural heritage while enriching fields like linguistics and anthropology. However, these languages are typically low-resource, making their reconstruction labor-intensive and costly. This challenge is exemplified by Nüshu, a rare script historically used by Yao women in China for self-expression within a patriarchal society. To address this challenge, we introduce NüshuRescue, an AI-driven framework designed to train large language models (LLMs) on endangered languages with minimal data. NüshuRescue automates evaluation and expands target corpora to accelerate linguistic revitalization. As a foundational component, we developed NCGold, a 500-sentence Nüshu-Chinese parallel corpus, the first publicly available dataset of its kind. Leveraging GPT-4-Turbo, with no prior exposure to Nüshu and only 35 short examples from NCGold, NüshuRescue achieved 48.69% translation accuracy on 50 withheld sentences and generated NCSilver, a set of 98 newly translated modern Chinese sentences of varying lengths. A sample of both NCGold and NCSilver is included in the Supplementary Materials. Additionally, we developed FastText-based and Seq2Seq models to further support research on Nüshu. NüshuRescue provides a versatile and scalable tool for the revitalization of endangered languages, minimizing the need for extensive human input.
@@
Ivory Yang1  Weicheng Ma2  Soroush Vosoughi3 1,2,3Department of Computer Science, Dartmouth College {Ivory.Yang.GR, Weicheng.Ma.GR, Soroush.Vosoughi}@dartmouth.edu
Figure 1: A simplified, stylized rendition of the NüshuRescue generation framework, using samples from NCGold
@@ 1 Introduction
Nüshu is a unique syllabic script originating from the ethnic Yao women in Jiangyong County, Hunan Province, China Broussard (2008); Liu (2014). This rare and historically significant script, often referred to as "women’s script" Wang (2020), is the only written language in the world developed by and used exclusively by women Ann-Gee (2013). In a past era of high female illiteracy, where women’s voices were largely suppressed Leung (2003), Nüshu served as a private means of communication and form of female empowerment in a male-dominated society Liu (1997). Today, Nüshu is an endangered language with very limited surviving documentation Luo et al. (2022), making it a critical focus for linguistic preservation and revitalization efforts Liu (2018).
@@
In the broader field of Natural Language Processing (NLP), low-resource languages like Nüshu present significant challenges, such as the scarcity of available data Magueresse et al. (2020) and absence of established linguistic tools Joshi et al. (2019); Krasadakis et al. (2024). The preservation of such languages is not only vital for maintaining linguistic diversity Gibbs (2002); Brenzinger and de Graaf (2009), but also for safeguarding the cultural heritage Romaine (2007) and historical narratives Hale (1992) they encapsulate. To preserve the Nüshu language, this paper provides two main contributions: (1) NüshuRescue: An LLM-based data generation framework to construct or scale up text corpora for low-resource languages, using minimal human labor and without presuming any prior knowledge of the target languages, and (2) Datasets: NCGold and NCSilver. NCGold is an expert-validated parallel corpus of 500 Nüshu-Chinese translation sentence pairs, while NCSilver is a Nüshu corpus generated through the NüshuRescue framework. Both NCGold and NCSilver are the first publicly available datasets of their kind.
Regarding language resources in Nüshu,  Zhao (1992) presented “A Compendium of Chinese Nüshu”, the most comprehensive known compilation of Nüshu scripts and their corresponding formal Chinese translations [^1] The very first version of this book was published in 1992; we referenced the most recent e-book released in 2019., featuring scans of original calligraphy artwork found since the 1950s Yan (2023). A sample is shown in Figure A1 in the Appendix. As part of our original contributions, we meticulously transcribed these paper-and-ink manuscripts into Unicode format for both Nüshu and Chinese, carefully categorizing and selecting key sentences to create the first digital dataset of its kind: NCGold. Given the limited material, we developed NüshuRescue, a data generation framework requiring only a small set of expert-validated ground-truth data. Using the NCGold dataset, it teaches an LLM to understand Nüshu and generate additional translations. Details of the manual annotation and generation framework are provided in Section 4, and summarized in Figure 1.
By leveraging NüshuRescue and a select 85 samples from NCGold (35 for instruction-tuning, 50 for evaluation), we translated 98 Chinese sentences of varying lengths from the UD Chinese GSDSimp dataset Wang et al. (2021) to Nüshu, forming the NCSilver dataset. Our framework is domain- and dataset-independent, allowing for easy scaling of Nüshu corpora. We then combined NCGold and NCSilver to form a total of 598 Nüshu-Chinese sentence pairs, and used this joint corpus to train a FastText model Athiwaratkun et al. (2018). Subsequently, we trained a Seq2Seq model Sriram et al. (2017) with the FastText word embeddings to (1) validate the quality of the NCGold and NCSilver datasets and (2) produce a clear-box, interpretable Machine Translation (MT) model between Nüshu and Chinese. Our results indicate that as data size increases, model performance improves, with BLEU, METEOR, and ROUGE scores rising from near 0 to 0.1365, 0.0942, and 0.1609, respectively, as shown in Table 2. This underscores the importance of data volume for training FastText and MT models, and highlights the value of the NüshuRescue framework in the study of endangered languages like Nüshu.
@@
Through our experiments, we find that modern AI models like GPT-4-Turbo demonstrate strong language learning capabilities, even in languages with limited prior exposure, allowing the NüshuRescue framework to effectively generate high-quality text for truly resource-scarce languages. While human expertise remains crucial for annotating seed data and ensuring translation accuracy, NüshuRescue significantly reduces the amount of human labor required compared to fully manual data annotation. Therefore, we believe that our NüshuRescue framework offers a promising solution for preserving endangered languages.
@@ 2 Related Work
Nüshu has garnered renewed cultural interest in recent years Thurnell-Read et al. (2022), especially after receiving international representation in the Hollywood production, "Snow Flower and the Secret Fan" Dargis (2011). However, existing linguistic work on Nüshu is limited; "A Compendium of Chinese Nüshu" Zhao (1992) remains the only formal print publication with expert-validated translations. Chiu (2012) provided a new perspective in "Heroines of Jiangyong" by compiling Nüshu scripts alongside English translations of their Chinese counterparts. While this is one step removed from a direct Nüshu-to-English translation, and lacks the formal mapping of linguistic structures, it is nonetheless valuable for introducing Nüshu to the English-speaking domain and expanding its accessibility.
Within the NLP community, there have been only tentative attempts to explore Nüshu. Tang et al. (2023) and Sun et al. (2023) proposed a two-part project combining research and art, where two AI agents trained on a Chinese-Nüshu dictionary collaborate to create a standard writing system for encoding Chinese, blending linguistic analysis with artistic interpretation. However, no publicly available digital dataset for Nüshu exists, and the use of NLP methods to analyze its syntax and semantics remains virtually unexplored. The only significant digitalization effort to date has been the creation of a Nüshu-Chinese Unicode dictionary Contributors , which remains a work in progress as it continues to be expanded and standardized.
Though Nüshu is poised for further expansion and preservation efforts, efforts to document and preserve other under-resourced and endangered languages have been more thoroughly explored. A common challenge across endangered language preservation projects is the need to gather data at scale. One approach is the collection of text corpora through web crawling and data extraction, as demonstrated by de Jesus and Nunes (2024) for Tetun, a language of Timor-Leste. However, this method relies on the presence of online resources, which is not the case for Nüshu. Another approach, transcription and human annotation, is more costly and labor-intensive, as seen in efforts to preserve Yoruba, a West African language Fagbolu et al. (2015). Given these constraints, it is crucial to combine human annotation efforts with AI and NLP tools, so as to advance Nüshu preservation in a cost-effective and scalable manner.
@@ 3 The Nüshu Language
Nüshu is a writing system originally created and exclusively used by foot-bound women in rural China Bissessar (2013), and therefore possesses internal linguistic and literary rules difficult for those outside the community to fully grasp. Nüshu emerged independently as a phonetic writing system, borrowing sounds from spoken Chinese (particularly the local dialect) and representing them using a syllabic system Zhao (2014). Nüshu is syllabic, meaning it uses symbols to represent individual syllables in spoken language Silber (1995). In contrast, Chinese is logographic, meaning it uses characters that represent words or morphemes, with each character typically corresponding to a meaning rather than a specific sound Tan et al. (2001). For example, Figure 2 illustrates the Nüshu and Chinese characters for the word "ocean," shown on the left and right, respectively. In Chinese, the character for "ocean" is logographic and conveys the concept as a whole, with the "water radical" (the three strokes on the left of the word) indicating its connection to water. Removing the "water radical" changes the character’s meaning to "each" or "every", illustrating how radicals influence the meaning of Chinese characters. In contrast, the Nüshu symbol for "ocean" is syllabic, representing the phonetic sounds of the spoken word rather than the concept itself.
Figure 2: Comparison of mapped Chinese (logographic) and Nüshu (syllabic) characters for the word "ocean".
@@
The phonetic nature of Nüshu creates a multi-mapping challenge with Chinese, as shown in Figure 3. A single Nüshu symbol can correspond to multiple Chinese characters with the same pronunciation but different meanings (one-to-many mapping), while a single Chinese character may map to several Nüshu symbols representing different phonetic variations (many-to-one mapping) Moratto (2022). Although an estimated 600–700 Nüshu characters exist, only 398 have been officially encoded in Unicode Noto Sans Nüshu Emerson (2002), leaving many symbols undocumented and creating gaps in translation efforts. Traditional methods like word alignment models Och (2003) and statistical machine translation Sharma et al. (2023), which rely on one-to-one mappings and large, parallel corpora, struggle to handle Nüshu’s complex relationships and limited dataset. Advanced techniques are needed to accommodate Nüshu’s unique structure and support corpus expansion and alignment with other languages.
@@ 4 NüshuRescue
4.1 Manual Data Annotation
@@
"A Compendium of Chinese Nüshu" Zhao (1992) is the most comprehensive, expert-validated collection of Nüshu calligraphy available Lo (2013). It features scanned Nüshu scripts, created with paper and ink, positioned at the top of each page, with the corresponding Chinese translation at the bottom. A sample from the book is shown in Figure A1 in the Appendix. While most works consist of traditional ancient Chinese poems and literary pieces, the collection also includes personal diary entries and an impassioned soliloquy written in response to a mother’s scolding.
Figure 4: Manual Data Annotation: Step-by-step transcription process
@@
The manual data annotation process for the 500-sentence NCGold dataset was carried out with meticulous attention to detail. The annotators for the task are college-educated, bilingual in English and Chinese, and have experience in computational linguistics. First, each sentence of the Chinese text translations was manually typed out, with the punctuation left out. Next, the Chinese text was copied and pasted into the official online website, Converter of Calligraphy Copybook of Nüshu Standard Characters Converter , to retrieve all possible mapped Nüshu words for each Chinese word. Finally, the annotators carefully filtered and matched the single correct Nüshu word to each corresponding Chinese word. Of the 500 total sentences, 127 had full coverage in the official Nüshu Unicode dictionary, with every Chinese word included. This step-by-step process is illustrated in Figure 4. For the remaining 373 sentences, while some Chinese words were covered, others were missing from the limited vocabulary of the dictionary, meaning the online converter could not provide Nüshu mappings for all terms. In these ambiguous cases, we manually consulted the Nüshu column of the official dictionary and visually identified the correct Nüshu characters from the book. As a result, NCGold not only provides the first expert-validated Nüshu-Chinese parallel corpus, but also contributes newly mapped Nüshu-Chinese word pairs, a small sample of which is shown in Figure A2 in the Appendix.
4.2 LLM-Based Data Annotation
@@
After obtaining NCGold, we established a text translation framework for scaling up the dataset by leveraging LLMs. For clarity, we refer to the manually-translated dataset as NCGold, and the subsequent LLM-generated dataset as NCSilver. One notable merit of our approach is that the LLM does not require any prior knowledge of Nüshu characters before being introduced to the provided examples, as demonstrated in Figure 5. While we use the GPT-4-Turbo model in our experiments, our proposed framework is model-agnostic, allowing for the use of any LLM with comparable language modeling and reasoning capabilities, as illustrated in Figure A3 in the Appendix.
Figure 5: GPT-4-Turbo does not understand Nüshu at the beginning
Figure 6: Prompt used on GPT-4-Turbo
@@
4.2.1 Experimental Settings
In our framework for obtaining Nüshu translations from the LLMs, we first provide translation examples from NCGold and the official Nüshu-Chinese dictionary for GPT-4-Turbo to learn from, uploading them as files to avoid excessively long prompts. The model is then prompted to generate Nüshu translations for Chinese input sentences, as shown in Figure 6. For quality assurance, 35 NCGold samples were provided for learning, and a round of evaluations was conducted on 50 withheld samples. We manually reviewed the LLM-generated translations for validation, correcting minor character choice errors. The choice of 35 samples was informed by preliminary tests, where fewer examples often led the model to refuse responses. Since there is no strict upper limit for the number of NCGold examples (aside from token limits), incorporating more high-quality examples could further enhance translation performance.
@@ 4.2.1 Experimental Settings
The UD Chinese GSDSimp dataset Wang et al. (2021) was used to construct a set of new Chinese sentences for annotation by GPT-4-Turbo. First, we filtered out sentences containing out-of-dictionary words, based on our Nüshu-Chinese dictionary, and then categorized the remaining sentences by ascending length. Table 1 details the data samples annotated in each round, along with the average sentence length. Arabic numbers were also removed to avoid biasing the translation evaluations. Through iterative annotations using ChatGPT and manual corrections, we progressively incorporated longer sentences into the translation process, ultimately creating a more comprehensive set of Nüshu-Chinese translation instances. Since all NCGold sentences are short, with a maximum length of 18 characters, we considered the potential challenges that LLMs might face in effectively translating longer sentences.
Figure 7: Pipeline for LLM-in-the-loop data annotation approach (Rounds 1-6)
Figure 8: Translation success and failure
4.2.2 Annotation Pipeline Setup

### v4

@@ Abstract
The preservation and revitalization of endangered and extinct languages is a meaningful endeavor, conserving cultural heritage while enriching fields like linguistics and anthropology. However, these languages are typically low-resource, making their reconstruction labor-intensive and costly. This challenge is exemplified by Nüshu, a rare script historically used by Yao women in China for self-expression within a patriarchal society. To address this challenge, we introduce NüshuRescue, an AI-driven framework designed to train large language models (LLMs) on endangered languages with minimal data. NüshuRescue automates evaluation and expands target corpora to accelerate linguistic revitalization. As a foundational component, we developed NCGold, a 500-sentence Nüshu-Chinese parallel corpus, the first publicly available dataset of its kind. Leveraging GPT-4-Turbo, with no prior exposure to Nüshu and only 35 short examples from NCGold, NüshuRescue achieved 48.69% translation accuracy on 50 withheld sentences and generated NCSilver, a set of 98 newly translated modern Chinese sentences of varying lengths. In addition, we developed FastText-based and Seq2Seq models to further support research on Nüshu. NüshuRescue provides a versatile and scalable tool for the revitalization of endangered languages, minimizing the need for extensive human input. All datasets and code have been made publicly available at https://github.com/ivoryayang/NushuRescue.
@@
Ivory Yang1  Weicheng Ma2  Soroush Vosoughi3 1,2,3Department of Computer Science, Dartmouth College {Ivory.Yang.GR, Weicheng.Ma, Soroush.Vosoughi}@dartmouth.edu
Figure 1: A simplified, stylized rendition of the NüshuRescue generation framework, using samples from NCGold.
@@ 1 Introduction
Nüshu is a unique syllabic script originating from the ethnic Yao women in Jiangyong County, Hunan Province, China Broussard (2008); Liu (2014). This rare and historically significant script, often referred to as “women’s script” Wang (2020), is the only written language in the world developed by and used exclusively by women Ann-Gee (2013). In a past era of high female illiteracy, where women’s voices were largely suppressed Leung (2003), Nüshu served as a private means of communication and form of female empowerment in a male-dominated society Liu (1997). Today, Nüshu is an endangered language with very limited surviving documentation Luo et al. (2022), making it a critical focus for linguistic preservation and revitalization efforts Liu (2018).
@@
In the broader field of Natural Language Processing (NLP), low-resource languages like Nüshu present significant challenges, such as the scarcity of available data Magueresse et al. (2020) and the absence of established linguistic tools Joshi et al. (2019); Krasadakis et al. (2024). The preservation of such languages is not only vital for maintaining linguistic diversity Gibbs (2002); Brenzinger and de Graaf (2009), but also for safeguarding the cultural heritage Romaine (2007) and historical narratives Hale (1992) they encapsulate. To preserve the Nüshu language, this paper provides two main contributions: (1) NüshuRescue: A LLM-based data generation framework to construct or scale up text corpora for low-resource languages, using minimal human labor and without presuming any prior knowledge of the target languages. (2.1) NCGold: An expert-validated parallel corpus of 500 Nüshu-Chinese translation sentence pairs. (2.2) NCSilver: A parallel corpus of 98 Nüshu-Chinese translation sentence pairs generated through the NüshuRescue framework. Both NCGold and NCSilver are the first publicly available datasets of their kind.
@@
Through our experiments, we find that LLMs like GPT-4-Turbo demonstrate strong language learning capabilities, even for languages with limited prior exposure, allowing the NüshuRescue framework to effectively generate higher-quality text for truly resource-scarce languages. While human expertise remains crucial for annotating seed data and ensuring translation accuracy, NüshuRescue significantly reduces the amount of human labor required compared to fully manual data annotation. As such, we believe that our NüshuRescue framework, as shown in Figure 1, offers a promising solution for the preservation and revitalization of endangered languages.
@@ 2 Related Work
Nüshu has garnered renewed cultural interest in recent years Thurnell-Read et al. (2022), especially after receiving international representation in the Hollywood production “Snow Flower and the Secret Fan” Dargis (2011). However, existing linguistic work on Nüshu is limited; “A Compendium of Chinese Nüshu” Zhao (1992) [^1] The very first version of this book was published in 1992; we referenced the most recent e-book released in 2019. remains the only formal print publication with expert-validated translations. Chiu (2012) provided a new perspective in “Heroines of Jiangyong” by compiling Nüshu scripts alongside English translations of their Chinese counterparts. While this is one step removed from a direct Nüshu-to-English translation and lacks formal linguistic mapping, it is nonetheless valuable for introducing Nüshu to the English-speaking domain.
There have been only tentative attempts to explore Nüshu within the NLP community. Tang et al. (2023) and Sun et al. (2023) proposed a two-part project combining research and art, where two AI agents trained on a Chinese-Nüshu dictionary collaborate to create a standard writing system for encoding Chinese, blending linguistic analysis with artistic interpretation. However, no publicly available digital dataset for Nüshu exists, and using NLP methods to analyze its syntax and semantics remains virtually unexplored. The only significant digitalization effort to date has been the creation of a Nüshu-Chinese Unicode dictionary Contributors , which remains a work in progress as it continues to be expanded and standardized.
Figure 2: Comparison of mapped Chinese (logographic) and Nüshu (syllabic) characters for the word “ocean”.
Efforts to document and preserve other under-resourced and endangered languages have been more thoroughly explored. A common challenge across endangered language preservation projects is the need to gather data at scale. One approach is the collection of text corpora through web crawling and data extraction, as demonstrated by de Jesus and Nunes (2024) for Tetun, a language of Timor-Leste. However, this method relies on the presence of online resources, which is not the case for Nüshu. Another approach, transcription, and human annotation is more costly and labor-intensive, as seen in efforts to preserve Yoruba, a West African language Fagbolu et al. (2015). Given these constraints, it is crucial to combine human annotation efforts with AI and NLP tools, so as to advance Nüshu preservation in a cost-effective and scalable manner.
@@ 3 The Nüshu Language
Nüshu is a writing system originally created and exclusively used by ethnic Yao women in rural China Bissessar (2013), and therefore possesses internal linguistic and literary rules difficult for those outside the community to fully grasp. Nüshu emerged independently as a phonetic writing system, borrowing sounds from spoken Chinese (particularly the local dialect) and representing them using a syllabic system Zhao (2014). Nüshu is syllabic, meaning it uses symbols to represent individual syllables in spoken language Silber (1995). In contrast, Chinese is logographic, meaning it uses characters that represent words or morphemes, with each character typically corresponding to a meaning rather than a specific sound Tan et al. (2001). For example, Figure 2 illustrates the Nüshu and Chinese characters for the word “ocean”, shown on the left and right, respectively. In Chinese, the character for “ocean” is logographic and conveys the concept as a whole, with the “water radical” (the three strokes on the left of the word) indicating its connection to water. Removing the “water radical” changes the character’s meaning to “each” or “every”, illustrating how radicals [^2] In Chinese, radicals are the building blocks of Chinese characters, often used as a means to organize them in dictionaries. They usually represent common semantic or phonetic components. influence the meaning of Chinese characters. In contrast, the Nüshu symbol for “ocean” is syllabic, representing the phonetic sounds of the spoken word rather than the concept itself.
@@
The phonetic nature of Nüshu creates a multi-mapping challenge with Chinese, as shown in Figure 3. A single Nüshu symbol can correspond to multiple Chinese characters with the same pronunciation but different meanings (one-to-many mapping). In contrast, a single Chinese character may map to several Nüshu symbols representing different phonetic variations (many-to-one mapping) Moratto (2022). Although an estimated 600–700 Nüshu characters exist, only 398 have been officially encoded in Unicode Noto Sans Nüshu Emerson (2002), leaving many symbols undocumented and creating gaps in translation efforts. Traditional methods like word alignment models Och (2003) and statistical machine translation Sharma et al. (2023), which rely on large, parallel corpora, struggle to handle Nüshu’s complex relationships and limited dataset. Advanced techniques are needed to accommodate Nüshu’s unique structure and support corpus expansion and alignment with other languages.
@@ 4 NüshuRescue
4.1 Manual Data Collection
@@
"A Compendium of Chinese Nüshu" Zhao (1992) is the most comprehensive, expert-validated collection of Nüshu calligraphy available Lo (2013). It features scanned Nüshu scripts, created with paper and ink, positioned at the top of each page, with the corresponding Chinese translation at the bottom. A sample from the book is shown in Figure A1 in the Appendix. While most works consist of traditional ancient Chinese poems and literature, the collection also includes personal diary entries and an impassioned soliloquy written in response to a mother’s scolding, providing contextual diversity.
Figure 4: Manual Data Collection: Step-by-step transcription process.
@@
The manual data collection process for the 500-sentence NCGold dataset was carried out with meticulous attention to detail. The annotators for the task are college-educated, bilingual in English and Chinese, and have experience in computational linguistics. The step-by-step process is illustrated in Figure 4333English Translation: “A noble man empowers the beauty within others”.. First, each sentence of the Chinese text translations was manually typed out, with the punctuation left out. Next, the Chinese text was copied and pasted into the official online website, Converter of Calligraphy Copybook of Nüshu Standard Characters Converter , to retrieve all possible mapped Nüshu words for each Chinese word. Finally, the annotators carefully filtered and matched the single correct Nüshu word to each corresponding Chinese word.
Figure 5: A small sample of newly mapped words not in the official Nüshu-Chinese dictionary; an original contribution.
Of the 500 total sentences, 127 had full coverage in the official Nüshu Unicode dictionary, with every Chinese word included. For the remaining 373 sentences, while some Chinese words were covered, others were missing from the limited vocabulary of the dictionary, meaning the online converter could not provide Nüshu mappings for all terms. In these ambiguous cases, we manually consulted the Nüshu column of the official dictionary and visually identified the correct Nüshu characters from the book. As a result, NCGold not only provides the first expert-validated Nüshu-Chinese parallel corpus but also contributes newly mapped Nüshu-Chinese word pairs, a small sample of which is shown in Figure 5 (Figure A2 in the Appendix shows a higher resolution version).
4.2 LLM-Based Data Augmentation
After obtaining NCGold, we established a text translation framework for scaling up the dataset by leveraging LLMs. For clarity, we refer to the manually-translated dataset as NCGold and the subsequent LLM-generated dataset as NCSilver.
@@ 4.2 LLM-Based Data Augmentation
One notable merit of our approach is that the LLM does not require any prior knowledge of Nüshu characters before being introduced to the provided examples, as demonstrated in Figure 6  [^4] English Translation: “A fragrant plum blossom comes from the bitter winter cold”. This is a traditional Chinese idiom that means for good things to happen, one often needs to go through hardship. While we use the GPT-4-Turbo model in our experiments, our proposed framework is model-agnostic, allowing for the use of any LLM with comparable language modeling and reasoning capabilities, as illustrated in Figure A3 in the Appendix.
Figure 6: GPT-4-Turbo does not understand Nüshu at the beginning.
Figure 7: Prompt used on GPT-4-Turbo.
4.2.1 Experimental Settings
In our framework for obtaining Nüshu translations from the LLMs, we first provide translation examples from NCGold and the official Nüshu-Chinese dictionary for GPT-4-Turbo to learn from, uploading them as files to avoid excessively long prompts. The model is then prompted to generate Nüshu translations for Chinese input sentences, as shown in Figure 7.
For quality assurance, 35 NCGold samples were provided for learning, and a round of evaluations was conducted on 50 withheld samples. We manually reviewed the LLM-generated translations for validation, correcting minor character choice errors. The choice of 35 samples was informed by preliminary tests, where fewer examples often led the model to refuse responses. Since there is no strict upper limit for the number of NCGold examples (aside from token limits), incorporating more high-quality examples could further enhance translation performance.
@@ 4.2.1 Experimental Settings
The UD Chinese GSDSimp dataset Wang et al. (2021) was used to construct a set of new Chinese sentences for annotation by GPT-4-Turbo. First, we filtered out sentences containing out-of-dictionary words based on our Nüshu-Chinese dictionary, and then categorized the remaining sentences by ascending length. Table 1 details the data samples annotated in each round, along with the average sentence length. Arabic numbers were also removed to avoid biasing the translation evaluations.
Through iterative annotations using ChatGPT and manual corrections, we progressively expanded the translation process to include longer sentences, resulting in a more comprehensive set of Nüshu-Chinese translation instances. This iterative approach to sentence length enabled the generation of progressively longer translations, surpassing the limitations of the existing dataset. However, translation accuracy for longer sentences was slightly lower than for shorter ones, likely due to the fact that all seed sentences were derived from short, ancient Chinese poems.
4.2.2 Generation Pipeline Setup